# Probabilistic Kernel Function for Fast Angle Testing

### v1

@@ 3.1 Reference-Angle-Based Design
(1)If $\cos\theta\leq\cos\phi_{1}=\left\langle\bm{q},\bm{v}\right\rangle$, then $\mathbb{P}[K(\bm{q},\bm{v})\geq\cos\theta]\geq p_{1}(\phi_{1})$
(2)If $\left\langle\bm{q},\bm{v}\right\rangle=\cos\phi_{2}<\cos\theta$, then $\mathbb{P}[K(\bm{q},\bm{v})\geq\cos\theta]<p_{2}(\phi_{2})$
, where $p_{2}(\phi_{2})$ is a strictly decreasing function in $\phi_{2}$ and $p_{1}(\phi_{1})>p_{2}(\phi_{2})$ when $\phi_{1}<\phi_{2}$.
@@
$F_{K^{1}_{S}(\bm{q},\bm{v})\mid A_{S}(\bm{q})}(x\mid\cos\psi)=I_{t}(\frac{d-2}{2},\frac{d-2}{2})$ (4)
@@
(2)Let $\bm{q}$, $\bm{v_{1}}$ and $\bm{v_{2}}$ be three normalized vectors on $\mathbb{S}^{d-1}$ such that $\left\langle\bm{q},\bm{v_{1}}\right\rangle>\left\langle\bm{q},\bm{v_{2}}\right\rangle$. The probability $\mathbb{P}[K^{1}_{S}(\bm{q},\bm{v_{1}})>K^{1}_{S}(\bm{q},\bm{v_{2}})|A_{S}(\bm{q})=\cos\psi]$ increases as $\psi$ decreases in $(0,\pi)$. In particular, when $\psi\in(0,\pi/2)$, $P[K^{1}_{S}(\bm{q},\bm{v_{1}})>K^{1}_{S}(\bm{q},\bm{v_{2}})|A_{S}(\bm{q})=\cos\psi]>0.5$.
Lemma 3.3
Let $\psi\in(0,\pi/2)$, that is, $A_{S}(\bm{v})=\cos\psi\in(0,1)$, and $d\geq 3$. Then $K^{2}_{S}$ is an angle-sensitive function, where $p_{1}(\phi)=0.5$ and $p_{2}(\phi)=I_{t^{\prime}}(\frac{d-2}{2},\frac{d-2}{2})$, where $t^{\prime}=\frac{1}{2}-\frac{\cos\theta-\cos\phi}{2\sin\phi\tan\psi}$.
@@
(2) (Deterministic relationship for angle testing) Lemma 3.2 establishes a relationship between the objective angle $\phi$ and the value of the function $K^{1}_{S}$. Notably, after computing $Z_{S}(\cdot)$, the value of reference angle $A_{S}(\cdot)$ can be obtained automatically. Besides, as will be shown in Sec. 3.2, with a reasonable choice of $S$, the assumption $A_{S}(\cdot)>0$ can always be ensured. Hence, eq. (4) essentially describes a deterministic relationship. In contrast to the asymptotic relationship of CEOs, eq. (4) provides an exact relationship without additional assumptions.
(3) (Effectiveness of kernel functions) The above two lemmas show that, with a reasonable construction of $S$ such that the reference angle is small with high probability, $K^{1}_{S}$ and $K^{2}_{S}$ can effectively address the corresponding angle testing problems. Specifically, the smaller the reference angle is, the more effective $K^{1}_{S}$ and $K^{2}_{S}$ become (By $p_{2}(\phi)$, a smaller $\psi$ leads to a smaller $p_{2}(\phi)$).
@@
3.2 Configuration of Points Set on Hypersphere
The remaining task is to configure the set $S$. Given $m$, our goal is to construct a set $S$ of $m$ points on $\mathbb{S}^{d-1}$, denoted by $S_{m}$, such that the reference angle $A_{S_{m}}(\cdot)$ is minimized. Due to the effect of the random rotation matrix $H$, the optimal configurations denoted by $\bar{S}_{m}$ and ${S}_{m}^{*}$, can be obtained either in the sense of expectation or in the sense of the worst case, respectively. That is,
$\bar{S}_{m}=\mathop{\rm argmax}_{S=\{u_{1},\dots,u_{m}\}\subset\mathbb{S}^{d-1}}\{{\mathbb{E}_{v\in U(\mathbb{S}^{d})}[A_{S}(\bm{v})]}\}.$ (5)
${S}_{m}^{*}=\underset{S=\begin{subarray}{c}\{u_{1},\dots,u_{m}\}\subset\mathbb{S}^{d-1}\end{subarray}}{\arg\max}\;\min_{v\in\mathbb{S}^{d-1}}\max_{1\leq i\leq m}\langle\bm{u_{i}},\bm{v}\rangle.$ (6)
By the definitions of $\bar{S}_{m}$ and ${S}_{m}^{\ast}$, they correspond to the configurations that achieve the smallest expected value and the smallest maximum value of $A_{S}(\bm{v})$, respectively. On the other hand, finding the exact solutions for $\bar{S}_{m}$ and ${S}_{m}^{\ast}$ is closely related to the best covering problem on the sphere, which is highly challenging and remains open in the general case. To the best of the authors’ knowledge, the optimal configuration ${S}_{m}^{\ast}$ is only known when $m\leq d+3$. In light of this, we provide two configurations of $S$: one relies on random antipodal projections (Alg. 1), and the other is built using multiple cross-polytopes (Alg. 2). Each has its own advantages. Specifically, Alg. 1 enables the estimation of reference angles, while Alg. 2 can empirically produce slightly smaller reference angles and is more efficient for projection computation.
@@ 3.2 Configuration of Points Set on Hypersphere
5             Generate a random rotation matrix $H\in\mathbb{R}^{d^{\prime}\times d^{\prime}}$, and rotate $2d$ axes in $\mathbb{R}^{d^{\prime}}$ using $H$  Collect the $2d$ vectors of the cross-polytope after rotation
@@ 7             Repeat the above iteration and select $b/2$ antipodal pairs from the rotated cross-polytope
$J(S)=\mathbb{E}_{v\in U(\mathbb{S}^{d})}[(A_{S}(\bm{v}))].$ (7)
By definition, $J(S)$ denotes the expected value of the cosine of the reference angles w.r.t $S$. This quantity is consistent with our theory, as a random rotation is applied to $\bm{v}$ or $\bm{q}$ in eq. (2) and eq. (3). Based on the previous analysis, for a fixed $m$, $J(S)$ is minimized when $S=\bar{S}_{m}$, which is hard to compute. However, we have the following result to approximately evaluate $J(S)$.
Lemma 3.4
[…]
Lemma 3.4 shows that, when $N$ is sufficiently large, we can approximately evaluate the performance of different configurations $S$’s by comparing their $\tilde{J}(S,N)$’s.
@@
(1) (Utilization of antipodal pairs and cross-polytopes) We use the antipodal pair or the cross-polytope as our building block for the following three reasons. (1) Since all the projection vectors are antipodal pairs, the evaluation time of projection vectors can be halved. (2) Both of two structures can ensure that the assumption $A_{S}(\bm{v})>0$ holds, such that the condition in Lemma 3.3 is always satisfied. (3) The result in antipodal-result shows that, for $m=2d$, under mild conditions, the $2d$ vertices of a cross-polytope can be proven to have the smallest covering radius, that is, the smallest reference angle in the worst case. Although the results in the case $m>2d$ are unknown, we can rotate the fixed cross-polytope in random directions to generate multiple cross-polytopes until we obtain $m$ vectors, which explains the steps from 6 to 10 in Alg. 2.
@@
(2) (Selection from random configurations) We can generate such $m$ points in the above way many times, which forms multiple $S$’s. By Lemma 3.4, we can use $\tilde{J}(S,N)$ to approximately evaluate the performance of $J(S)$, and thus, among the generated $S$’s, we select the configuration $S_{\text{pol}}$ corresponding to the maximal $\tilde{J}(S,n)$. This explains steps 5 and 11 in Alg. 2.
@@
(3) (Accuracy boosting via multiple levels) Clearly, increasing $m$ can lead to a smaller reference angle. The analysis in peos shows that, for certain angle-thresholding problems requiring high accuracy, an exponential increase in $m$, rather than a linear one, can be effective. Therefore, similar to peos , we use a product-quantization-like technique PQ to partition the original space into $L$ subspaces (levels), which is adopted in both algorithms. By concatenating equal-length sub-vectors from these $L$ subspaces, we can virtually generate $m^{L}$ normalized projected vectors. As will be shown in Lemma 3.5, the introduction of $L$ can significantly decrease the reference angle.
@@
(4) (Fast projection computation via multi-cross-polytypes) In eq. (2) and eq. (3), we need to compute $H\bm{q}$, $H\bm{v}$, $HS\bm{q}$ and $HS\bm{v}$ in the indexing phase or in the query phase. If such projection time is a concern in practice, we can set $R$ and $L$ to 1 in Alg. 2 to accelerate the projection without the explicit computation of $S$. Specifically, if $m\geq 2d$, suppose that $m$ is divisible by $2d$ and $m=2ld$. $H^{\top}$ or $(HS)^{\top}$ can be approximated by $[S_{(1)},S_{(2)},S_{(l)}]^{\top}$, where $S_{(i)}$ here denotes the $i$-th structured random projection matrix. Then the projection time cost for $v$ or $q$ can be reduced from $O(md)$ to $O((m\log d)/2)$. If $2d>m$, the cost is $O(d\log d)$ by the projection matrix completion.
By Alg. 1 and Alg. 2, we obtain structures $S_{\text{sym}}(m,L)$ and $S_{\text{pol}}(m,L)$ virtually containing $m^{L}$ projection vectors. For $S_{\text{sym}}(m,L)$, we can establish a relationship between $J(S_{\text{sym}}(m,L))$ and $(m,L)$ as follows, which reveals the impact of the choice of $(m,L)$ on reference angle.
Lemma 3.5
[…]
The RHS of ineq. (8) actually denotes $J(S_{\text{ran}})$, where $S_{\text{ran}}$ is the configuration of purely random projections (see Alg. 3 and Sec. A.4 in Appendix for more detail). A numerical computation of the RHS of ineq. (8) is shown in the Appendix (Fig. 4). On the other hand, due to the introduction of cross-polytopes, the analysis of $J(S_{\text{pol}}(m,L))$ is challenging. However, simulation experiments show that $J(S_{\text{pol}}(m,L))$ is only slightly larger than $J(S_{\text{sym}}(m,L))$, which makes the lower bound in ineq. (8) still applicable to $J(S_{\text{pol}}(m,L))$ in practice.
@@ 4.1 Improvement on CEOs-Based Techniques
As for $K^{1}_{S}$, we can use it to improve CEOs, which is used for Maximum Inner Product Search (MIPS) and further applied to accelerate LSH-based ANNS falconn and DBSCAN sDBSCAN . Since CEOs is originally designed for inner products, we generalize $K^{1}_{S}$ to $K^{1^{\prime}}_{S}$ as follows to align with CEOs:
@@
It is easy to see that, with two minor modifications, that is, replacing $\bm{v}\in\mathbb{S}^{d-1}$ with $\bm{v}\in\mathbb{R}^{d}$, and replacing $x$ with $\|\bm{v}\|x$ in eq. (4), Lemma 3.2 still holds. Therefore, $K^{1^{\prime}}_{S}$ can be regarded as a reasonable kernel function for inner products. Then, we can apply $K^{1^{\prime}}_{S}$ to the algorithm in ceos ; Falconn++ ; sDBSCAN . We only need to make the following modification. In these algorithms, the random Gaussian matrix, which denotes the set of projection vectors, can be replaced by $S_{\text{sym}}$ or $S_{\text{pol}}$, with the other parts unchanged. This substitution does not change the complexity of the original algorithms. To distinguish this projection technique based on $K^{1^{\prime}}_{S}$ from CEOs, we refer to it as KS1 (see Alg. 4 for the projection structure of KS1). In the experiments, we will demonstrate that KS1 yields a slight improvement in recall rates over CEOs, owing to a smaller reference angle.

### v2

@@ 4 Analysis of Probability Guarantees
(1) If $\cos\theta\leq\cos\phi_{1}=\left\langle\bm{q},\bm{v}\right\rangle$, then $\mathbb{P}[K(\bm{q},\bm{v})\geq\cos\theta]\geq p_{1}(\phi_{1})$,
(2) If $\left\langle\bm{q},\bm{v}\right\rangle=\cos\phi_{2}<\cos\theta$, then $\mathbb{P}[K(\bm{q},\bm{v})\geq\cos\theta]<p_{2}(\phi_{2})$,
where $p_{2}(\phi_{2})$ is a strictly decreasing function in $\phi_{2}$ and $p_{1}(\phi_{1})>p_{2}(\phi_{2})$ when $\phi_{1}<\phi_{2}$.
@@
$F_{K^{1}_{S}(\bm{q},\bm{v})}(x\mid A_{S}(\bm{q})=\cos\psi)=I_{t}\left(\frac{d-2}{2},\frac{d-2}{2}\right),$ (4)
@@
(2) Let $\bm{q}$, $\bm{v_{1}}$ and $\bm{v_{2}}$ be three normalized vectors on $\mathbb{S}^{d-1}$ such that $\left\langle\bm{q},\bm{v_{1}}\right\rangle>\left\langle\bm{q},\bm{v_{2}}\right\rangle$. The probability $\mathbb{P}[K^{1}_{S}(\bm{q},\bm{v_{1}})>K^{1}_{S}(\bm{q},\bm{v_{2}})|A_{S}(\bm{q})=\cos\psi]$ increases as $\psi$ decreases in $(0,\pi)$. In particular, when $\psi\in(0,\pi/2)$, $P[K^{1}_{S}(\bm{q},\bm{v_{1}})>K^{1}_{S}(\bm{q},\bm{v_{2}})|A_{S}(\bm{q})=\cos\psi]>0.5$, that is, $K^{1}_{S}$ satisfies the probability guarantee in Problem 1.1.
Lemma 4.3.
Let $\psi\in(0,\pi/2)$, that is, $A_{S}(\bm{v})=\cos\psi\in(0,1)$, and $d\geq 3$. Then $K^{2}_{S}$ is an angle-sensitive function. Precisely, $K^{2}_{S}$ satisfies the probability guarantee in Problem 1.2, where $\epsilon_{1}=0.5$ and $\epsilon_{2}=I_{t^{\prime}}(\frac{d-2}{2},\frac{d-2}{2})<0.5$, where $t^{\prime}=\frac{1}{2}-\frac{\cos\theta-\cos\phi}{2\sin\phi\tan\psi}$.
@@
(2) (Deterministic relationship for angle testing) Lemma 4.2 establishes a relationship between the objective angle $\phi$ and the value of the function $K^{1}_{S}$. Notably, after computing $Z_{S}(\cdot)$, the value of reference angle $A_{S}(\cdot)$ can be obtained automatically. Besides, as will be shown in Sec. 5, with a reasonable choice of $S$, the assumption $A_{S}(\cdot)>0$ can always be ensured. Hence, Eq. (4) essentially describes a deterministic relationship. In contrast to the asymptotic relationship of CEOs, Eq. (4) provides an exact relationship without additional assumptions.
(3) (Effectiveness of kernel functions) The above two lemmas show that, with a reasonable construction of $S$ such that the reference angle is small with a high probability, $K^{1}_{S}$ and $K^{2}_{S}$ can effectively address the corresponding angle testing problems. The smaller the reference angle is, the more effective $K^{1}_{S}$ and $K^{2}_{S}$ become.
@@
5 Implementation and Complexity Analysis
We discuss how to configure $S$, and then analyze the complexity of $K^{1}_{S}$ and $K^{2}_{S}$. Based on the discussion in Sec. 4, we observe that small reference angles are preferred. Thus, given $m$, our goal is to construct a set $S$ of $m$ points on $\mathbb{S}^{d-1}$, denoted by $S_{m}$, such that the reference angle $A_{S_{m}}(\cdot)$ is minimized. Due to the effect of the random rotation matrix $H$, the optimal configurations denoted by $\bar{S}_{m}$ and ${S}_{m}^{*}$, can be obtained either in the sense of expectation or in the sense of the worst case, respectively:
$\bar{S}_{m}=\mathop{\rm argmax}_{S=\{\bm{u_{1}},\dots,\bm{u_{m}}\}\subset\mathbb{S}^{d-1}}\{{\mathbb{E}_{\bm{v}\in U(\mathbb{S}^{d-1})}[A_{S}(\bm{v})]}\},$ (5)
${S}_{m}^{*}=\underset{S=\begin{subarray}{c}\{\bm{u_{1}},\dots,\bm{u_{m}}\}\subset\mathbb{S}^{d-1}\end{subarray}}{\arg\max}\;\min_{\bm{v}\in\mathbb{S}^{d-1}}\max_{1\leq i\leq m}\langle\bm{u_{i}},\bm{v}\rangle,$ (6)
where $U(\mathbb{S}^{d-1})$ denotes the uniform distribution on the sphere. By the definitions of $\bar{S}_{m}$ and ${S}_{m}^{\ast}$, they correspond to the configurations that achieve the smallest expected value and the smallest maximum value of $A_{S}(\bm{v})$, respectively. On the other hand, finding the exact solutions for $\bar{S}_{m}$ and ${S}_{m}^{\ast}$ is closely related to the best covering problem, which is highly challenging and remains open in the general case. To the best of the authors’ knowledge, the optimal configuration ${S}_{m}^{\ast}$ is only known when $m\leq d+3$ (Borodachov et al., 2019). In light of this, we provide two configurations of $S$: one relies on random antipodal projections (Alg. 1), and the other is built using multiple cross-polytopes (Alg. 2). Each has its own advantages. Alg. 1 enables the estimation of reference angles, while Alg. 2 can empirically produce slightly smaller reference angles and is more efficient for projection computation.
@@ 5 Implementation and Complexity Analysis
3      Generate a random rotation matrix $H\in\mathbb{R}^{d^{\prime}\times d^{\prime}}$, and rotate $2d^{\prime}$ axes in $\mathbb{R}^{d^{\prime}}$ using $H$  Collect the $2d^{\prime}$ vectors of the cross-polytope after rotation
@@ 5      Repeat the above iteration and select $b/2$ antipodal pairs from the rotated cross-polytope
$J(S)=\mathbb{E}_{\bm{v}\in U(\mathbb{S}^{d-1})}[(A_{S}(\bm{v}))].$ (7)
By definition, $J(S)$ denotes the expected value of the cosine of the reference angles w.r.t $S$. This quantity is consistent with our theory, as a random rotation is applied to $\bm{v}$ or $\bm{q}$ in Eq. (2) and Eq. (3). Based on the previous discussion, for a fixed $m$, $J(S)$ is minimized when $S=\bar{S}_{m}$, which is hard to compute. Thus, we take $N$ to be sufficiently large and let $v_{1,N},\ldots,v_{N,N}$ be vectors drawn independently and uniformly from $U(\mathbb{S}^{d-1})$. Let $\tilde{J}(S,N)=[\Sigma_{i=1}^{N}A_{S}(\bm{v_{i,N}})]/N$. By the law of large number, we can approximate $J(S)$ by $\tilde{J}(S,N)$ when $N$ is sufficiently large.
@@
(1) (Utilization of antipodal pairs and cross-polytopes) We use the antipodal pair or the cross-polytope as our building block for the following three reasons. (i) Since all the projection vectors are antipodal pairs, the evaluation time of projection vectors can be halved. (ii) Both of two structures can ensure that the assumption $A_{S}(\bm{v})>0$ holds, such that the condition in Lemma 4.3 is always satisfied. (iii) The result in Borodachov (2022) shows that, for $m=2d$, under mild conditions, the $2d$ vertices of a cross-polytope can be proven to have the smallest covering radius, that is, the smallest reference angle in the worst case. Although the results in the case $m>2d$ are unknown, we can rotate the fixed cross-polytope in random directions to generate multiple cross-polytopes until we obtain $m$ vectors, which explains the steps from 3 to 7 in Alg. 2.
@@
(2) (Selection from random configurations) We can generate such $m$ points in the above way many times, which forms multiple $S$’s. By the discussion above, we can use $\tilde{J}(S,N)$ to approximately evaluate the performance of $J(S)$, and thus, among the generated $S$’s, we select the configuration $S_{\text{pol}}$ corresponding to the maximal $\tilde{J}(S,N)$. This explains steps 2 and 8 in Alg. 2.
@@
(3) (Accuracy boosting via multiple levels) Increasing $m$ can lead to a smaller reference angle. The analysis in Lu et al. (2024) shows that, for certain angle-thresholding problems requiring high accuracy, an exponential increase in $m$, rather than a linear one, can be effective. Therefore, similar to Lu et al. (2024), we use a product-quantization-like technique (Jégou et al., 2011) to partition the original space into $L$ subspaces (levels), which is adopted in both algorithms. By concatenating equal-length sub-vectors from these $L$ subspaces, we can virtually generate $m^{L}$ normalized projected vectors. As will be shown in Lemma B.1, the introduction of $L$ can significantly decrease the reference angle.
@@
(4) (Fast projection computation via multi-cross-polytypes) In Eq. (2) and Eq. (3), we need to compute $HS\bm{v}$ in the indexing phase. If the projection time is a concern in practice, we can use the Fast Johnson–Lindenstrauss transformation to accelerate this process. Specifically, we use Alg. 2 with $R=1$. When $L=1$, the cost of computing $HS\bm{v}$ is $O(\max(d,m)\log d)$. When $L>1$, the cost of computing $HS\bm{v}$ can be reduced to $O(d\log d+mL\log(d/L))$.
By Alg. 1 and Alg. 2, we obtain structures $S_{\text{sym}}(m,L)$ and $S_{\text{pol}}(m,L)$ virtually containing $m^{L}$ projection vectors. For $S_{\text{sym}}(m,L)$, we can actually establish a deterministic relationship between $J(S_{\text{sym}}(m,L))$ and $(m,L)$ (see Lemma B.1 in Appendix). On the other hand, in practice, $J(S_{\text{pol}}(m,L))$ is often slightly larger than $J(S_{\text{sym}}(m,L))$, as will be shown in Table 1.
With $S_{\text{sym}}(m,L)$ and $S_{\text{pol}}(m,L)$, we are ready to present a complexity analysis of the two proposed functions, showing that they satisfy the complexity requirement in Problem 1.1 and Problem 1.2.
Lemma 5.1.
In the indexing phase, for a fixed dataset $\mathcal{D}$ of size $n$, the complexities of $K^{1}_{S}$ and $K^{2}_{S}$ are $O(nmd)$ and $O(nd\log d+nmd)$, respectively. In the query phase, $K^{1}_{S}$ requires $O(md)$- time for random projection, and for each $v\in\mathcal{D}$, it spends $O(1)$-time for computing $K^{1}_{S}(q,v)$, while $K^{2}_{S}$ spends $O(d\log d+md)$-time for random projection and random rotation, and for each $v\in D$, it spends $O(L)$-time for computing $K^{2}_{S}(q,v)$, where $L\ll d$. Particularly, if $S=S_{\text{pol}}(m,L)$ with $R=1$, the complexities of $K^{1}_{S}$ and $K^{2}_{S}$ in the indexing phase can be reduced to $O(\max(d,m)n\log d)$ and $O(nd\log d+nmL\log(d/L))$, respectively, where $m$ is greater than $d/L$ by user-specification.
@@ 6.1 Improvement on CEOs-Based Techniques
As for $K^{1}_{S}$, we can use it to improve CEOs, which is used for MIPS and further applied to accelerate LSH-based ANNS (Andoni et al., 2015) and DBSCAN (Xu & Pham, 2024). Since CEOs is originally designed for inner products, we generalize $K^{1}_{S}$ to $K^{1^{\prime}}_{S}$ as follows to align with CEOs:
@@
It is easy to see that, with two minor modifications, that is, replacing $\bm{v}\in\mathbb{S}^{d-1}$ with $\bm{v}\in\mathbb{R}^{d}$, and replacing $x$ with $\|\bm{v}\|x$ in Eq. (4), Lemma 4.2 still holds. Therefore, $K^{1^{\prime}}_{S}$ can be regarded as a reasonable kernel function for inner products. Then, we can apply $K^{1^{\prime}}_{S}$ to the algorithm in Pham (2021); Pham & Liu (2022); Xu & Pham (2024). We only need to make the following modification. In these algorithms, the random Gaussian matrix, which denotes the set of projection vectors, can be replaced by $S_{\text{sym}}$ or $S_{\text{pol}}$, with the other parts unchanged. This substitution does not change the complexity of the original algorithms. To distinguish this projection technique based on $K^{1^{\prime}}_{S}$ from CEOs, we refer to it as KS1 (see Alg. 4 for the projection structure of KS1). In the experiments, we will demonstrate that KS1 yields a slight improvement in recall rates over CEOs, owing to a smaller reference angle.